MICROWAVE SPECTRA OF THE 2,4-PENTADIYNE RADICAL, H_{2}C-C≡C-C≡C-H

WEI CHEN, Harvard-Smithsonian Center for Astrophysics, Cambridge, MA 02138; and Division of Engineering and Applied Sciences, Harvard University, Cambridge, MA 02138; and Department of Chemistry, Wesleyan University, Middletown, CT 06459; \underline{STEWART E. NOVICK}, Department of Chemistry, Wesleyan University, Middletown, CT 06459; M. C. McCARTHY, P. THADDEUS, Harvard-Smithsonian Center for Astrophysics, Cambridge, MA 02138; and Division of Engineering and Applied Sciences, Harvard University, Cambridge, MA 02138; C. A. GOTTLIEB, Division of Engineering and Applied Sciences, Harvard University, Cambridge, MA 02138.

The H_{2}C-C≡C-C≡C-H radical was produced in a pulsed discharge supersonic jet and studied by Fourier transform microwave spectroscopy. Four rotational transitions of the radical from N_{upper} = 2 to 5 in the K_{a} = 0 ladder were observed between 8 and 21 GHz. A total of 36 hyperfine components were measured and assigned. Seven spectroscopic constants, including rotational and centrifugal distortion constants, the spin-rotation constant, and hyperfine coupling constants for the methylenic and acetylenic hydrogens, were determined. These spectroscopic constants of 2,4-pentadiyne will be compared with those of its isoelectronic analog, H_{2}C-C≡C-C≡N, and with those of the propargyl radical, H_{2}C-C≡C-H.